My question is "Use the definition of a parabola and the distance formula to find the equation of a parabola with directix y=2 and focus (-3,6)"
Please help! Also looking for a possible online math tutor. Please e-mail me with information, Thanks!

Answer by scott8148(3751)

"definition of a parabola" __ the locus of points equidistant from a point (focus) and a line (directrix)

for a point (x,y) __ the distance from y=2 is y-2

the distance from (-3,6) is sqrt((x+3)^2+(y-6)^2)

(y-2)^2=(x+3)^2+(y-6)^2 __ y^2-4y+4=x^2+6x+9+y^2-12y+36 __ 8y=x^2+6x+41

y=(x^2+6x+41)/8